# OlympiadBench: A Challenging Benchmark for Promoting AGI with Olympiad-Level Bilingual Multimodal Scientific Problems

### v1

@@ Abstract
Recent advancements have seen Large Language Models (LLMs) and Large Multimodal Models (LMMs) surpassing general human capabilities in various tasks, approaching the proficiency level of human experts across multiple domains. With traditional benchmarks becoming less challenging for these models, new rigorous challenges are essential to gauge their advanced abilities. In this work, we present OlympiadBench, an Olympiad-level bilingual multimodal scientific benchmark, featuring 8,952 problems from Olympiad-level mathematics and physics competitions, including the Chinese college entrance exam. Each problem is detailed with expert-level annotations for step-by-step reasoning. Evaluating top-tier models on OlympiadBench, we implement a comprehensive assessment methodology to accurately evaluate model responses. Notably, the best-performing model, GPT-4V, attains an average score of 17.23% on OlympiadBench, with a mere 11.28% in physics, highlighting the benchmark rigor and the intricacy of physical reasoning. Our analysis orienting GPT-4V points out prevalent issues with hallucinations, knowledge omissions, and logical fallacies. We hope that our challenging benchmark can serve as a valuable resource for helping future AGI research endeavors. The data and evaluation code are available at https://github.com/OpenBMB/OlympiadBench
@@
[…]
@@ 1 Introduction
To address the aforementioned inadequacies, we introduce OlympiadBench, a Olympiad-level bilingual multimodal scientific benchmark. This collection comprises 8,952 math and physics problems sourced from International Olympiads, Chinese Olympiads, and the most challenging segments of the Chinese College Entrance Exam (GaoKao). We download PDF data from official websites and utilize Mathpix [^1] https://mathpix.com/ for OCR parsing. We meticulously inspect, clean, and revise the data, and further adopt LLMs for deduplication. Finally, we annotate the data with crucial information such as answer types and subfields, yielding a dataset that is clean, accurate, and detailed. As shown in Figure 1, OlympiadBench features numerous distinct characteristics such as difficulty, free-form generation, expert-level solution annotation, detailed labeling of difficulty, wide-coverage of modality and language, etc. These features are summarized more clearly from Table 1.
We evaluate a wide variety of LLMs and LMMs on OlympiadBench. GPT-4V, a fusion of the strongest LLMs and LMMs, achieves a score of 20.35% in mathematics, 11.28% in physics. Importantly, the experiment results show that LMMs still struggle in computational error, incorrect reasoning or induction. For the process involved in the correct responses, the process occasionally includes hallucinated reasoning, or choosing a more complex solution when a simpler solution exists. All these results highlight the substantial challenge our benchmark presents to contemporary models and point the direction of future efforts.
@@ 2 Related Work
Physics Benchmarks. Physics questions in SciQ Welbl et al. (2017) and ScienceQA Lu et al. (2022) are mainly elementary and high school level multiple-choice questions, lacking complex reasoning and computational tasks. In MMLU-STEM Hendrycks et al. (2020) and C-Eval-STEM Huang et al. (2023), physics questions also adopt a multiple-choice format. JEEBench Arora et al. (2023) extends this format to include multistep reasoning with physics knowledge, yet it is limited in scope and purely text-only. SciEval Sun et al. (2023) consists of a total of about 18,000 challenging scientific questions, spanning three important basic science fields: chemistry, physics and biology. SciBench Wang et al. (2024) and OCWCourses Lewkowycz et al. (2022) offer college-level physics questions in free-response formats, where SciBench contains multimodal information. In contrast, OlympiadBench escalates in difficulty, diversifies in question types, and surpasses in volume, setting a new benchmark for complexity and variety in the domain.
@@ 3 The OlympiadBench Dataset
To evaluate the reasoning abilities of LLMs and LMMs in mathematics and physics problems, we have created OlympiadBench, a bilingual and multimodal scientific benchmark at the competition level. This section provides a detailed account of the construction process of OlympiadBench. Summarized statistics of the dataset is shown in Table 2, and more detailed statistics per subject are in Appendix A.2.
@@
Total Problems 8,952
   * Problems with images 5,129 (57%)
   * Problems with solutions 8,952 (100%)
@@
EN: ZH 2,288: 6,664
Open-ended Questions 7,254 (81%)
Theorem Proving 1,698 (19%)
Math: Physics 6,524: 2,428
   * Maths with images 3,102
   * Physics with images 2,027
@@
Max question tokens 3,701
Average solution tokens 352
Max solution tokens 4,213
@@ 3.1 Design Principle
The motivation behind the design of OlympiadBench is to establish a benchmark that represents the pinnacle of human intellectual achievement, thereby encouraging researchers of LLMs to push the boundaries of mathematical and physical reasoning capabilities. To realize this vision, we focus on curating challenges that epitomize the highest level of competition worldwide. Specifically, our benchmark includes:
@@
    Inclusion of Olympiad-Level Problems. The chosen competition problems are of Olympiad caliber, aimed at the most accomplished students in a region’s high school education phase. These problems are open-ended, setting them apart from the conventional multiple-choice or fill-in-the-blank formats. This selection is designed to more accurately capture the complexity of advanced mathematical reasoning.
@@
    Provision of Detailed Solutions. Given the advanced difficulty of these problems, which may exceed the comprehension of individuals without a specialized background in mathematics, each problem is accompanied by expertly crafted solutions that detail the reasoning steps involved. This approach can not only reduces the difficulty of annotation and evaluation but also enhances the accuracy of the solutions provided.
@@
    Incorporation of Visuals. Recognizing the crucial role of visual information in conveying complex ideas, our benchmark incorporates problems that require understanding images, identifying spatial relationships, and other advanced reasoning tasks. This inclusion aims to assess and improve the model’s capabilities in interpreting visual data as part of its reasoning process.
@@
Through these meticulously planned criteria, OlympiadBench aspires to not only challenge but also to significantly advance the frontier of LLM capabilities in mathematical and physical reasoning.
@@ 3.2 Data Processing
Data Collection. OlympiadBench is meticulously compiled from three primary sources: Global Mathematics and Physics Olympiad Problems, Regional and National Chinese Math Competitions, and Gaokao Mock Questions for Mathematics and Physics   [^3] Clean and correct college entrance exams as well as mock questions from Wisdom Way AI Lab.. Each chosen for its distinct advantages in creating a robust and comprehensive benchmark for evaluating LLMs and LMMs in mathematical and scientific reasoning. Their challenges progressively increase in difficulty, not only distinguishing the reasoning capabilities of models of various sizes but also offering guidance on scaling laws for specialized models in these domains.
Format Conversion and Deduplication. After collecting all PDF files, we utilize the Mathpix  [^4] https://mathpix.com/ tool for OCR recognition and convert them into markdown format. However, no conversion process is flawless, necessitating manual verification by our team members between the original PDF files and the converted Markdown texts. The Markdown texts are further structured into a format akin to "Problem—Solution—Answer", employing its markup language for text organization. Subsequently, we leverage a specialized small-scale language model  [^5] https://huggingface.co/Laurie/Bloom1b7-deepspeed-chat-Chinese-math trained on mathematical symbol corpora for vectorizing the data and performing deduplication based on cosine similarity.
@@ 3.3 Data Characteristics
Progressive Problems in Physics. In physics competitions such as the International Physics Olympiad (IPhO), problems are often structured around a common material or scenario, with subsequent questions potentially relying on the answers or information from previous questions. One example is given in Figure 7 and Figure 8 This design characteristic is commonly referred to as "Progressive Problems." By linking a series of questions together, progressive problems require participants to apply their knowledge and skills comprehensively to gradually solve more complex issues. This type of question design aims to test students’ depth of understanding, application capabilities, and innovative thinking, rather than just basic knowledge. To better utilize this feature, we have compiled the material, questions, and their answers that precede each progressive problem into its "Context" field.
Answer Type Classification. Whether in mathematics or physics, the answers to problems requiring definitive responses can largely be categorized into the following types: numeric, expression, equation, interval, and tuple. Simple examples of these can be seen in Table 3.
@@ 4.1 Settings
We conduct evaluations of open-source and closed-sourced LMMs that have been selected with consideration of their comprehensive capabilities on OlympiadBench. At the same time, we have selected LLMs with strong mathematical and logical abilities for evaluation on plain text questions.
@@ 4.1.2 Evaluation Workflow
We first apply each model to generate answers for questions in OlympiadBench using prompts formed by prompt template, with open-ended models running on NVIDIA A800 GPUs. Then, we run the automatic scoring pipeline to judge the correctness of the answers as described in subsection 3.4. Finally, we calculate the micro-average accuracy as the comparing metric. The code of the whole workflow is provided in the supplementary material.
@@ 4.2 Baselines
In our study, we evaluate the performance of current leading bilingual large multimodal models (LMMs), as well as bilingual large language models (LLMs) that has strong mathematical and reasoning abilities. We take both open- and closed-source models into consideration, and use the largest and latest released checkpoint or the best-performing official API that can be achieved.
For LMMs, we selected GPT-4V(GPT-4-Vision) OpenAI (2023b), Gemini-Pro-Vision Team (2023), Qwen-VL-Max Bai et al. (2023) for closed-source models, while Yi-VL-34B 01-ai (2024) and LLaVA-NeXT-34B Liu et al. (2024) for open-source models. For models that demand compulsory image input, we take their LMM counterpart (corresponding text-model api or base LLM) for evaluation.To examine the impact of replacing LMM with base LLM for processing text-only data, we subsequently compare the performance differences between GPT-4V and GPT-4  [^7] The version of GPT-4 is ”0125-preview” and GPT-4V is ”1106-vision-preview”. on text-only questions in OlympiadBench.
@@ 4.3 Main Results
LLaVA-NEXT-34B†  [^8] Conducted text-only experiment using corresponding foundation LLM: Nous-Hermes-2-Yi-34B NousResearch (2023) 3.88 2.43 4.85 4.36 - 2.12 1.69 1.80 3.60
Yi-VL-34B†  [^9] Conducted text-only experiment using corresponding foundation LLM: Yi-34B-Chat 01-ai (2023) 4.38 5.11 4.68 4.67 - 0.95 1.75 1.55 3.37
@@
LLaVA-NEXT-34B88footnotemark: 8 3.64 2.71 9.88 6.39 - 2.16 5.45 1.98 5.80
Yi-VL-34B99footnotemark: 9 4.18 5.30 9.08 6.69 - 0.87 9.08 1.21 5.95
DeepSeekMath-7B-RL 19.73 2.71 27.54 20.25 - 6.49 15.45 9.38 18.73
@@
Qwen-VL-Max 11.68* 14.29 27.34 19.83 - 4.37* 21.82 10.03 18.47
GPT-4V 29.51* 16.01 38.31* 31.20 - 12.99 25.45 16.98 29.24
GPT-4 30.69* 16.50 38.48 31.75 - 11.44 28.18 16.84 29.50
Table 4: Experimental results. En_COMP: COMP problems in English, Zh_COMP: COMP problems in Chinese, Zh_CEE: CEE problems in Chinese. For closed-source models, the responses for some problems are not available, we mark the results with * (all of the proportion of missing answers are less than 5%). The causes are further described in Appendix B.3. Moreover, LLaVA-NEXT-34B and Yi-VL-34B only accepts input with single image, we mark results from only one image input with †.
@@
OlympiadBench is more challenging than existing benchmarks, which provides new perspective to compare LMMs. As shown in table 7, the most advanced model only achieves an average accuracy of 17.23% on OlympiadBench, which is much lower than that of existing benchmarks. Moreover, the gap between the models has been widened, thereby becoming more significant, which helps people to compare the differences in capabilities between different models more accurately.
@@ 5.2 Mistake Analysis of GPT-4V
We manually sample and check 97 maths (55 for English and 42 for Chinese) and 67 physics Olympics-level open-ended problems that GPT-4V fails, and analyze the type of mistakes, the overall results are shown in figure 4. In maths problems, the typical errors of GPT-4V include: insufficient classification discussion, especially in combinatorial problems; poor performance in problems requiring large calculations (e.g. conic curve problems), manifests as a lack of logic in the calculation process, resulting in the model being unable to provide a reasonable answer. However, we also found that GPT-4V has strong abilities in solving quadratic equations and derivative problems. In physics problems, GPT-4V tends to fall in conceptual confusion, or introduce unnecessary variables or concepts, but its capability to simplify and transform algebraic expressions is stronger than in purely mathematical situations, with nearly no numerical calculation errors.

### v2

@@ Abstract
Recent advancements have seen Large Language Models (LLMs) and Large Multimodal Models (LMMs) surpassing general human capabilities in various tasks, approaching the proficiency level of human experts across multiple domains. With traditional benchmarks becoming less challenging for these models, new rigorous challenges are essential to gauge their advanced abilities. In this work, we present OlympiadBench, an Olympiad-level bilingual multimodal scientific benchmark, featuring 8,476 problems from Olympiad-level mathematics and physics competitions, including the Chinese college entrance exam. Each problem is detailed with expert-level annotations for step-by-step reasoning. Evaluating top-tier models on OlympiadBench, we implement a comprehensive assessment methodology to accurately evaluate model responses. Notably, the best-performing model, GPT-4V, attains an average score of 17.97% on OlympiadBench, with a mere 10.74% in physics, highlighting the benchmark rigor and the intricacy of physical reasoning. Our analysis orienting GPT-4V points out prevalent issues with hallucinations, knowledge omissions, and logical fallacies. We hope that our challenging benchmark can serve as a valuable resource for helping future AGI research endeavors. The data and evaluation code are available at https://github.com/OpenBMB/OlympiadBench
@@
Chaoqun He1, Renjie Luo2, Yuzhuo Bai1, Shengding Hu1, Zhen Leng Thai1 Junhao Shen1, Jinyi Hu1, Xu Han1∗, Yujie Huang1, Yuxiang Zhang3 Jie Liu3, Lei Qi3, Zhiyuan Liu1thanks: Corresponding authors: Xu Han and Zhiyuan Liu., Maosong Sun1 1Dept. of Comp. Sci. & Tech., Institute for AI, Tsinghua University, Beijing, China 2Institute of Artificial Intelligence, Beihang University, China 3Wisdom Way AI Lab, China {hcq21,byz22}@mails.tsinghua.edu.cn, renjie.luo@outlook.com, {hanxu2022,liuzy}@tsinghua.edu.cn
@@ 1 Introduction
To address the aforementioned inadequacies, we introduce OlympiadBench, an Olympiad-level bilingual multimodal scientific benchmark. This collection comprises 8,476 math and physics problems sourced from International Olympiads, Chinese Olympiads, and the most challenging segments of the Chinese College Entrance Exam (GaoKao). We download PDF data from official websites and utilize Mathpix [^1] https://mathpix.com/ for OCR parsing. We meticulously inspect, clean, and revise the data, and further adopt LLMs for deduplication. Finally, we annotate the data with crucial information such as answer types and subfields, yielding a dataset that is clean, accurate, and detailed. As shown in Figure 1, OlympiadBench features numerous distinct characteristics such as difficulty, free-form generation, expert-level solution annotation, detailed labeling of difficulty, wide-coverage of modality and language, etc. These features are summarized more clearly from Table 1.
We conduct an evaluation of current state-of-the-art LLMs and LMMs on the OlympiadBench. The best-performing model, GPT-4V, is a multimodal version of GPT-4 developed by OpenAI that can understand images. Despite its advanced capabilities, GPT-4V achieves a score of only 17.97% on OlympiadBench, with individual scores of 21.70% in mathematics and 10.74% in physics.
Importantly, the experiment results show that LMMs still struggle in computational error, incorrect reasoning or induction. For the process involved in the correct responses, the process occasionally includes hallucinated reasoning, or choosing a more complex solution when a simpler solution exists. All these results highlight the substantial challenge OlympiadBench presents to contemporary large models and point the direction of future efforts.
@@ 2 Related Work
Physics Benchmarks. Physics questions in SciQ Welbl et al. (2017), ScienceQA Lu et al. (2022) and E-EVAL Hou et al. (2024) are mainly elementary and high school level multiple-choice questions, lacking complex reasoning and computational tasks. In MMLU-STEM Hendrycks et al. (2020) and C-Eval-STEM Huang et al. (2023), physics questions also adopt a multiple-choice format. JEEBench Arora et al. (2023) extends this format to include multistep reasoning with physics knowledge, yet it is limited in scope and purely text-only. TheoremQA Chen et al. (2023) is the first theorem-driven question-answering dataset. Curated by domain experts, it contains 800 high-quality questions that encompass 350 theorems from Mathematics, Physics, Electrical Engineering and Computer Science (EE&CS), and Finance. SciEval Sun et al. (2023) consists of a total of about 18,000 challenging scientific questions, spanning three important basic science fields: chemistry, physics and biology. SciBench Wang et al. (2024) and OCWCourses Lewkowycz et al. (2022) offer college-level physics questions in free-response formats, where SciBench contains multimodal information. In contrast, OlympiadBench escalates in difficulty, diversifies in question types, and surpasses in volume, setting a new benchmark for complexity and variety in the domain.
@@ 3 The OlympiadBench Dataset
To evaluate the reasoning abilities of LLMs and LMMs in mathematics and physics problems, we have created OlympiadBench, a bilingual and multimodal scientific benchmark at the competition level. This section provides a detailed account of the construction process of OlympiadBench. Summarized statistics of the dataset is shown in Table 2, and more detailed statistics are in Appendix A.2.
@@
Total Problems 8,476
* Problems with images 4,869 (57%)
* Problems with solutions 8,476 (100%)
@@
EN: ZH 2,125: 6,351
Open-ended Questions 6,728 (79%)
Theorem Proving 1,748 (21%)
Math: Physics 6,142: 2,334
* Maths with images 2,911
* Physics with images 1,958
@@
Max question tokens 3,745
Average solution tokens 347
Max solution tokens 4,223
@@ 3.1 Design Principle
The motivation behind OlympiadBench is to establish a benchmark that represents the pinnacle of human intellectual achievement, thereby encouraging researchers of large models to push the boundaries of mathematical and physical reasoning capabilities. We focus on curating challenges that epitomize the highest level of competition worldwide. Specifically, OlympiadBench includes:
@@
    Inclusion of Olympiad-Level Problems. We collect mathematics and physics problems from the International Olympiad competitions, which cater to the most outstanding high school students in a region. These problems are open-ended, differing from traditional multiple-choice or fill-in-the-blank formats. This selection aims to more accurately reflect the complexity of advanced scientific reasoning, providing insight into the actual reasoning process of the models.
@@
    Provision of Detailed Solutions. Given the advanced difficulty of these problems, which may exceed the comprehension of individuals without a specialized background in mathematics, each problem is accompanied by expertly crafted solutions that detail the reasoning steps involved. This approach can not only reduces the difficulty of annotation and evaluation but also enhances the accuracy of the solutions provided. Furthermore, detailed expert-level solutions are valuable for research in model reasoning.
@@
    Incorporation of Visuals. Recognizing the crucial role of visual information in conveying complex ideas, OlympiadBench incorporates problems that require understanding images, identifying spatial relationships, and other advanced reasoning tasks. This inclusion aims to assess the model’s capabilities in interpreting visual data as part of its reasoning process.
@@
Through these carefully designed criteria, OlympiadBench aims to not only challenge but also significantly advance the capabilities of models in mathematical and physical reasoning.
@@ 3.2 Data Processing
Data Collection. OlympiadBench is meticulously compiled from three primary sources: Global Mathematics and Physics Olympiad Problems, Regional and National Chinese Math Competitions, and Gaokao Mock Questions for Mathematics and Physics   [^3] Clean and correct college entrance exams as well as mock questions from Wisdom Way AI Lab.. Each chosen for its distinct advantages in creating a robust and comprehensive benchmark for evaluating LLMs and LMMs in mathematical and scientific reasoning. Their challenges progressively increase in difficulty, not only distinguishing the reasoning capabilities of models of various sizes but also offering guidance on scaling laws Kaplan et al. (2020) for specialized models in these domains.
Format Conversion and Deduplication. After collecting all PDF files, we utilize the Mathpix tool for OCR recognition and convert them into markdown format. However, no conversion process is flawless, necessitating manual verification by our team members between the original PDF files and the converted Markdown texts. The Markdown texts are further structured into a format akin to "Problem—Solution—Answer", employing its markup language for text organization. Subsequently, we leverage a specialized small-scale language model  [^4] https://huggingface.co/Laurie/Bloom1b7-deepspeed-chat-Chinese-math trained on mathematical symbol corpora for vectorizing the data and performing deduplication based on cosine similarity.
@@ 3.3 Data Characteristics
Progressive Problems in Physics. In physics competitions such as the International Physics Olympiad (IPhO), problems are often structured around a common material or scenario, with subsequent questions potentially relying on the answers or information from previous questions. One example is given in Figure 10 and Figure 11. This design characteristic is commonly referred to as "progressive problems". By linking a series of questions together, progressive problems require participants to apply their knowledge and skills comprehensively to gradually solve more complex issues. This type of question design aims to test students’ depth of understanding, application capabilities, and innovative thinking, rather than just basic knowledge. To better utilize this feature, we compile the material, questions, and their answers into the ’context’ field for each set of progressive problems.
Answer Type Classification. Whether in mathematics or physics, the answers to problem requiring definitive response can largely be categorized into the following types: numeric, expression, equation, interval, and tuple. Simple examples of these can be seen in Table 3.
@@ 4.1 Settings
We conduct evaluations of open-source and closed-sourced LMMs that have been selected with consideration of their comprehensive capabilities on OlympiadBench. At the same time, we have selected LLMs with strong mathematical and logical abilities for evaluation on text-only questions.
@@ 4.1.2 Evaluation Workflow
We first apply each model to generate answers for questions in OlympiadBench using prompts formed by prompt template, with open-source models running on NVIDIA A800 GPUs. Then, we run the automatic scoring pipeline to judge the correctness of the answers as described in subsection 3.4. Finally, we calculate the micro-average accuracy as the comparing metric. The code of the whole workflow is provided in the supplementary material.
@@ 4.2 Baselines
In our study, we evaluate the performance of current leading bilingual large multimodal models (LMMs), as well as bilingual large language models (LLMs) that has strong mathematical and reasoning abilities. We take both open- and closed-source models into consideration, using either the largest and latest released checkpoints or the best-performing official APIs available.
For LMMs, we select GPT-4V(GPT-4-Vision) OpenAI (2023b), Gemini-Pro-Vision Team (2023), Qwen-VL-Max Bai et al. (2023) for closed-source models, while Yi-VL-34B 01-ai (2024) and LLaVA-NeXT-34B Liu et al. (2024) for open-source models. For models that demand compulsory image input, we take their LMM counterpart (corresponding text-model API or base LLM) for evaluation. Specifically, for LLaVA-NeXT-34B, we use its base LLM, Nous-Hermes-2-Yi-34B NousResearch (2023). The text model corresponding to Yi-VL-34B is Yi-34B-Chat 01-ai (2023). Similarly, for the Gemini-Pro-Vision, we utilize the Gemini-Pro API interface. To examine the impact of replacing LMM with base LLM for processing text-only data, we subsequently compare the performance differences between GPT-4V and GPT-4  [^6] The version of GPT-4 is ”0125-preview” and GPT-4V is ”1106-vision-preview”. on text-only questions in OlympiadBench.
@@ 4.3 Main Results
LLaVA-NeXT-34B† 3.98 2.60 4.64 4.30 - 1.36 2.32 2.08 3.65
Yi-VL-34B† 4.22 3.68 4.31 4.23 - 0.91 1.64 1.46 3.42
@@
LLaVA-NeXT-34B 4.15 2.94 8.55 6.29 - 2.12 5.22 3.13 5.87
Yi-VL-34B 4.45 3.68 8.06 6.24 - 0.85 5.22 2.28 5.72
DeepSeekMath-7B-RL 19.44 2.70 22.42 18.09 - 6.78 16.52 9.97 17.02
@@
Qwen-VL-Max 11.57 14.29 25.89 19.70 - 4.24 18.26 8.83 18.27
GPT-4V 28.93 15.93 37.10 31.01 - 12.71 23.48 16.24 29.07
GPT-4 30.42 16.42 37.98 32.00 - 12.29 24.35 16.24 29.93
Table 4: Experimental results. En_COMP: COMP problems in English, Zh_COMP: COMP problems in Chinese, Zh_CEE: CEE problems in Chinese. For closed-source models, the responses for some problems are not available, we mark the results with * (all of the proportion of missing answers are less than 1%, except for the result of Qwen-VL-Max in Physics-En_COMP, where 26 questions exceed maximum input length). The causes are further described in Appendix B.3. Moreover, LLaVA-NeXT-34B and Yi-VL-34B only accepts input with single image, we mark results from only one image input with †.
@@
OlympiadBench is more challenging than existing benchmarks, which provides new perspective to compare LMMs. As shown in table 7, the most advanced model only achieves an average accuracy of 17.97% on OlympiadBench, which is much lower than that of existing benchmarks. Moreover, the gap between the models has been widened, thereby becoming more significant, which helps people to compare the differences in capabilities between different models more accurately.
@@ 5.2 Mistake Analysis of GPT-4V
We manually sample and check 97 maths (55 for English and 42 for Chinese) and 67 physics Olympiad-level open-ended problems that GPT-4V fails, and analyze the type of mistakes, the overall results are shown in figure 4. In maths problems, the typical errors of GPT-4V include: insufficient classification discussion, especially in combinatorial problems; poor performance in problems requiring large calculations (e.g. conic curve problems), manifests as a lack of logic in the calculation process, resulting in the model being unable to provide a reasonable answer. However, we also found that GPT-4V has strong abilities in solving quadratic equations and derivative problems. In physics problems, GPT-4V tends to fall in conceptual confusion, or introduce unnecessary variables or concepts, but its capability to simplify and transform algebraic expressions is stronger than in purely mathematical situations, with nearly no numerical calculation errors.
6 Discussion and Future Work
Here, we discuss the automated evaluation of theorem problems and the future research directions in advanced disciplines.
Evaluation of Theorem’s Proof Problems. The inability to automatically evaluate theorem’s proof problems remains a significant challenge today. Currently, mainstream methods for the automatic evaluation of proofs require formalization, necessitating domain expertise and background knowledge. Exploring how to automatically evaluate natural language proofs represents an important research direction. Our dataset includes expert-level comprehensive annotations in a fusion of natural language and LaTeX formats, making it a high-quality resource for research in natural language proof problems and fostering further development.
Expansion of Disciplines. Mathematics, much like physics, serves as a litmus test for artificial intelligence, requiring a strong foundational knowledge, rigorous high-level computations, and precise logical reasoning. Currently, large models still face significant challenges in mathematics and physics, which are critical hurdles that must be overcome for the development of AGI. In our future work, we will integrate additional fields such as geography, biology, and chemistry to provide a more innovative and comprehensive evaluation of a model’s reasoning capabilities.